How many grams of magnesium do you need to produce 285 grams of magnesium chloride.
Mg + 2HCl----> MgCl2 + H2
1
Expert's answer
2013-05-22T11:30:24-0400
First of all we can find the amount of substance of magnesium chloride. We can use the formula
n = m/M
where
n - the amount of substance
m - the mass
M - the molar mass, which is calculated: M(MgCl2) = Ar(Mg) + 2* (Cl) = 24 + 2*35.5 = 95g/mol
n(MgCl2) = 285 g / 95 g/mol = 3 moles
According the equation Mg + 2HCl--->MgCl2 + H2 we can find the amount of substance of magnesium:
n(Mg) = n(MgCl2) = 3 moles
So the mass will be:
m(Mg) = n(Mg)*M(Mg)
M(Mg) = 24 g/mol
m(Mg) = 3 moles * 24 g/mol = 72 g

Answer: 72 g